# FLIGAN: Enhancing Federated Learning with Incomplete Data using GAN

### v1

@@ Abstract.
Federated Learning (FL) provides a privacy-preserving mechanism for distributed training of machine learning models on networked devices (e.g., mobile devices, IoT edge nodes). It enables Artificial Intelligence (AI) at the edge by creating models without sharing the actual data across the network. Existing research works typically focus on generic aspects of non-IID data and heterogeneity in client’s system characteristics, but they often neglect the issue of insufficient data for model development, which can arise from uneven class label distribution and highly variable data volumes across edge nodes. In this work, we propose FLIGAN, a novel approach to address the issue of data incompleteness in FL. First, we leverage Generative Adversarial Networks (GANs) to adeptly capture complex data distributions and generate synthetic data that closely resemble the real-world data. Then, we use synthetic data to enhance the robustness and completeness of datasets across nodes. Our methodology adheres to FL’s privacy requirements by generating synthetic data in a federated manner without sharing the actual data in the process. We incorporate techniques such as classwise sampling and node grouping, designed to improve the federated GAN’s performance, enabling the creation of high-quality synthetic datasets and facilitating efficient FL training. Empirical results from our experiments demonstrate that FLIGAN significantly improves the model accuracy, especially in scenarios with high class imbalances, achieving up to a 20% increase in model accuracy over traditional FL baselines.
Federated Learning, Incomplete Data, GANs, Edge Computing, Edge AI
@@ 1. Introduction
One of the main impeding challenges of FL is data incompleteness  (zhao2018federated, ), which can be defined with two scenarios: (i) Certain classes predominate on some nodes while being absent on others, and/or (ii) Data volume varies across nodes, with some possessing more data than others. The problem of data incompleteness is exacerbated in edge environments, as edge nodes often operate in unreliable conditions and can experience various issues such as temporal device failures, sensor malfunctioning, and network connectivity issues  (mcmahan2017communication, ; yang_2019_federated, ).
@@
While GANs have traditionally been applied successfully in the area of image and text synthesis (chang2020synthetic, ; mcmahan2017learning, ), our work explores their capability to generate tabular data within an FL framework in a privacy-preserving manner. Directly applying existing federated GAN models for decentralized tabular data presents some challenges. A significant issue is the multimodal distributions in continuous columns across nodes, leading to mode collapse and widening the discrepancy between synthetic and real data (xu2019modeling, ). Thus, in this work, we propose FLIGAN (Federated Learning with Incomplete data using GAN), which utilizes GANs in federated settings and addresses these challenges by adopting several techniques including classwise sampling and node grouping by intelligently clustering nodes based on their data quantity. This technique trains generators separately for each class label, thereby narrowing the gap between synthetic and real data in the process, as observed from the results. Furthermore, we use these generators in runtime and add synthetic data step by step in the training process of the FL classification model. Our experimental study on real-world datasets demonstrates that FLIGAN outperforms baselines by improving model accuracy.
@@ 5       Model $\leftarrow$ TrainFederatedModel(CombinedData)
Dataset Rows Cat. Cols Cont. Cols Total
@@ 6. Conclusions and Future Work
In this study, we introduce FLIGAN, an efficient method using federated GANs to synthetically augment incomplete data for improved model accuracy. We benchmarked FLIGAN against FedAvg and FedGAN across various datasets and showed improved accuracy, particularly with complex datasets like Intrusion and Adult. Although FedAvg is faster, FLIGAN’s accuracy gains demonstrate its effectiveness in FL environments with incomplete data. In the future, we plan to refine FLIGAN for greater accuracy and broader dataset compatibility, exploring advanced GAN architectures for different data types.
@@ Acknowledgements
This work has been partially funded through following projects: Transprecise Edge Computing (Triton), Austrian Science Fund (FWF): P 36870-N, Trustworthy and Sustainable Code Offloading (Themis), Austrian Science Fund (FWF): PAT1668223, and Satellite-based Monitoring of Livestock in the Alpine Region (Virtual Shepherd), FFG Austrian Space Applications Programme ASAP 2022 #53079251.

### v2

@@ Abstract.
Federated Learning (FL) provides a privacy-preserving mechanism for distributed training of machine learning models on networked devices (e.g., mobile devices, IoT edge nodes). It enables Artificial Intelligence (AI) at the edge by creating models without sharing actual data across the network. Existing research typically focuses on generic aspects of non-IID data and heterogeneity in client’s system characteristics, but they often neglect the issue of insufficient data for model development, which can arise from uneven class label distribution and highly variable data volumes across edge nodes. In this work, we propose FLIGAN, a novel approach to address the issue of data incompleteness in FL. First, we leverage Generative Adversarial Networks (GANs) to adeptly capture complex data distributions and generate synthetic data that closely resemble real-world data. Then, we use synthetic data to enhance the robustness and completeness of datasets across nodes. Our methodology adheres to FL’s privacy requirements by generating synthetic data in a federated manner without sharing the actual data in the process. We incorporate techniques such as classwise sampling and node grouping, designed to improve the federated GAN’s performance, enabling the creation of high-quality synthetic datasets and facilitating efficient FL training. Empirical results from our experiments demonstrate that FLIGAN significantly improves model accuracy, especially in scenarios with high class imbalances, achieving up to a 20% increase in model accuracy over traditional FL baselines.
journalyear: 2024copyright: rightsretainedconference: 7th International Workshop on Edge Systems, Analytics and Networking; April 22, 2024; Athens, Greecebooktitle: 7th International Workshop on Edge Systems, Analytics and Networking (EdgeSys ’24), April 22, 2024, Athens, Greece
@@ 1. Introduction
One of the main challenges of FL is data incompleteness  (zhao2018federated, ), which can be defined with two scenarios: (i) Certain classes predominate on some nodes while absent on others, and/or (ii) Data volume varies across nodes, with some possessing more data than others. The problem of data incompleteness is exacerbated in edge environments, as edge nodes often operate in unreliable conditions and can experience various issues such as temporal device failures, sensor malfunctioning, and network connectivity issues  (mcmahan2017communication, ; yang_2019_federated, ).
@@
While GANs have traditionally been applied in the area of image and text synthesis (chang2020synthetic, ; mcmahan2017learning, ), our work explores their capability to generate tabular data within an FL framework in a privacy-preserving manner. Directly applying existing federated GAN models for decentralized tabular data presents some challenges. A significant issue is the multimodal distributions in continuous columns across nodes, leading to mode collapse and widening the discrepancy between synthetic and real data (xu2019modeling, ). Thus, in this work, we propose FLIGAN (Federated Learning with Incomplete data using GAN), which utilizes GANs in federated settings and addresses these challenges by adopting several techniques including classwise sampling and node grouping by clustering nodes based on their data quantity. This technique trains generators separately for each class label, thereby narrowing the gap between synthetic and real data in the process, as observed from the results. Furthermore, we use these generators in runtime and add synthetic data step by step in the training process of the FL classification model. Our experimental study on real-world datasets demonstrates that FLIGAN outperforms baselines by improving model accuracy.
@@ 5       Model $\leftarrow$ TrainFederatedModel(CombinedData)
4. Performance Evaluation
4.1. Experimental setup
Data Rows Cat Cont Tot
@@ 6. Conclusions and Future Work
In this study, we introduce FLIGAN, an efficient method using federated GANs to synthetically augment incomplete data for improved accuracy. We benchmarked FLIGAN against FedAvg and FedGAN across various datasets and showed improved accuracy, particularly with complex datasets like Intrusion and Adult. Although FedAvg is faster, FLIGAN’s accuracy gains demonstrate its effectiveness in FL environments with incomplete data. In the future, we plan to refine FLIGAN for greater accuracy and broader compatibility, exploring advanced GAN architectures for different data types.
@@ Acknowledgements
This research was funded in part by the Austrian Science Fund (FWF) through following projects: Transprecise Edge Computing (Triton), Austrian Science Fund (FWF): 10.55776/P36870, Trustworthy and Sustainable Code Offloading (Themis), Austrian Science Fund (FWF): 10.55776/PAT1668223, and by the Austrian Research Promotion Agency (FFG) through the following project: Satellite-based Monitoring of Livestock in the Alpine Region (Virtual Shepherd), FFG Austrian Space Applications Programme ASAP 2022 #53079251.